# Bridging the Gap Between Information Seeking and Product Search Systems: Q&A Recommendation for E-commerce

### v1

@@ 1 Introduction
This process, however, is not optimal. First, users that search for products may not have enough knowledge to ask the right questions in an information seeking system. Second, an independent information seeking system does not directly take into account the full context of the user in product search (e.g., the product detail page that the user is viewing or the set of search results). For those reasons, as illustrated in Figure 1, users would often need to frequently switch between the two systems to refine their requirements and finally reach a purchase decision. This results in a lengthy process requiring substantial user effort to achieve their shopping goals. Some recent work has attempted to address these issues by enabling product search systems to answer informational queries [chen2023industry], but combining these distinct features in a single interface may not be intuitive for users.
@@ 4 Q&A Generation
For those reasons, we propose to use LLMs for the generation of Q&A pairs. Using LLMs is advantageous for the task since they can generate both the question and the answer from their internal knowledge [Hämäläinen et al., 2023; Hartvigsen et al., 2022]. To generate questions that are useful for the context, we can provide the context information as part of the input (e.g., the query, items in the SERP, or the product detail page). LLMs also have good capability to process long inputs, which means that we can provide it with different types of data to generate a variety of interesting questions. For example, for product detail page questions we can use customer reviews, the product specifications, and relevant external knowledge to generate questions. Some of these ideas have been explored by vedula2024question in the context of shopping assistants.
@@ 6 Optimizing Engagement with User Feedback Loops
How to optimize the Q&A recommendation system with LLMs is an important aspect of the system. While this topic has not been investigated extensively due the recency of LLM-based recommender systems, it is likely to attract growing interest. For example, senel2024generativeexploreexploittrainingfreeoptimization recently investigated how generative explore-exploit methods could be used to optimize cohort-level user engagement with LLM-generated content, as measured by click through rates (CTR). There are several challenges that need to be addressed for the effective optimization of the Q&A recommendation problem.

### v2

@@ 1 Introduction
This process, however, is not optimal. First, users that search for products may not have enough knowledge to ask the right questions in an information seeking system. Second, an independent information seeking system does not directly take into account the full context of the user in product search (e.g., the product detail page that the user is viewing or the set of search results). For those reasons, as illustrated in Figure 1, users would often need to frequently switch between the two systems to refine their requirements and finally reach a purchase decision. This results in a lengthy process requiring substantial user effort to achieve their shopping goals. Some recent work has attempted to address these issues by enabling product search systems to answer informational queries [Chen et al., 2023], but combining these distinct features in a single interface may not be intuitive for users.
@@ 4 Q&A Generation
For those reasons, we propose to use LLMs for the generation of Q&A pairs. Using LLMs is advantageous for the task since they can generate both the question and the answer from their internal knowledge [Hämäläinen et al., 2023; Hartvigsen et al., 2022]. To generate questions that are useful for the context, we can provide the context information as part of the input (e.g., the query, items in the SERP, or the product detail page). LLMs also have good capability to process long inputs, which means that we can provide it with different types of data to generate a variety of interesting questions. For example, for product detail page questions we can use customer reviews, the product specifications, and relevant external knowledge to generate questions. Some of these ideas have been explored by Vedula et al. [2024] in the context of shopping assistants.
@@ 6 Optimizing Engagement with User Feedback Loops
How to optimize the Q&A recommendation system with LLMs is an important aspect of the system. While this topic has not been investigated extensively due the recency of LLM-based recommender systems, it is likely to attract growing interest. For example, Senel et al. [2024] recently investigated how generative explore-exploit methods could be used to optimize cohort-level user engagement with LLM-generated content, as measured by click through rates (CTR). There are several challenges that need to be addressed for the effective optimization of the Q&A recommendation problem.